Mrs,Perez is 3 times as old as her son, jaypee in 12 years,Mrs Perez will be twice as old as Kaylee.
What is the present age. Thank you.

Answer by ikleyn(52790)

.
Mrs,Perez is 3 times as old as her son, Jaypee. In 12 years, Mrs. Perez will be twice as old as Kaylee. What is the present age. Thank you.
----------------------------------------------------------------

Let P = Mrs. Perez' present age,
    J = Jaypee      --- " -----.

Then you have two equations

P      = 3J,
P + 12 = 2(J+12).


Substitute (1) into (2). You will get

3J + 12 = 2(J+12  --->  3J + 12 = 2J + 24  ---> 3J - 2J = 24 - 12  --->  J = 12.

Answer. J = 12, P = 36.